Suppose country A produces 3 goods only; textbooks, computers and cell-phones. The total production and prices of all goods are given in the following table. Year 2000 is the base year.

Item

2000

2001

2002

Price

Quantity

Price

Quantity

Price

Quantity

Textbooks

$200

12

$250

16

$400

14

Computers

$1,000

5

$1,200

9

$1,500

7

Cell-phones

$600

10

$650

15

$800

12

  1. Calculate the nominal GDP in 2002.
  2. Calculate the real GDP in 2001 and 2002.
  3. Explain why real GDP decreased between 2001 and 2002.
  4. Without calculations, explain why nominal GDP increased between 2001 and 2002.
  5. Calculate the inflation rate between 2000 and 2002.
  6. Suppose students resold their textbooks. Will this affect GDP? Explain.
Answer #1

As per the HOMEWORKLIB POLICY, answering the first four parts

a.

Nominal GDP is calculated based on current year prices and quantities

Nominal GDP in 2002 = 400*14 + 1500*7 + 800*12

= 5,600 + 10,500 + 9,600

= 25,700

b

Real GDP is calculated based on current year quantities and base year prices.

Real GDP in 2001 = 200*16 + 1000*9 + 600*15

= 3200 + 9000 + 9000

= 21,200

Real GDP in 2002 = 200*14 + 1000*7 + 600*12

= 2800+ 7000 + 7200

= 17,000

c.

Real GDP takes the base year prices for its calculation. So, prices for 2000 year are fixed. We observed from the table that the quantity demanded/produced for Textbooks, Computers and Cell-phones have declined respectively.

This causes a fall in the Real GDP value in year 2002

d.

Nominal GDP in 2002 = 25,700

Nominal GDP in 2001 = 250*16 + 1200*9 + 650*15

= 4,000 + 10,800 + 9,750

= 24,550

Nominal GDP has increased from 24,550 in 2001 to 25,700 in 2002.

The value has increased because there has been an increase in the prices of all three goods in 2002 as compared to 2001. Since, Nominal GDP is not adjusted for inflation, it shows an inflated amount.